If sin inverse x + sin inverse y=π
and x=ky.... Find value of 39^2k + 5^k
 

Grade:12

1 Answers

Arun
5 years ago
sin^-1 x + sin^-1 y = pi/2 + pi/2
 
This is only possible if x = y = 1
 
Now
x = ky
1 = k *1 
 k = 1
 
Now
39^2 + 5^1 = 1521 + 5 = 1526